Logical Competence Versus Logical Correctness

Paula Bucholc

In the broad discussion devoted to logical competence almost nobody has addressed the question of the relationship between logical competence and logical correctness. A priori, it is not certain that logical competence is necessarily logically correct. Therefore, we have examined the important research of Johnson-Laird and Bara [Johnson–Laird, Bara 1984] from a new perspective. We were looking not for syllogistic conclusions that can be reached by people, but rather for the ability to recognise correct syllogistic reasoning.

Johnson-Laird and Bara treat some invalid syllogisms as valid, namely those in which their correctness require non-empty terms. Moreover the results of their investigations suggest that people are not ready to draw any existential conclusions from two universal premises, which is a case of any syllogism requiring for its correctness non-empty terms. It is known that human beings have had trouble with empty terms in syllogistic for more than 2000 years.

Therefore we have decided to look at the behaviour of human beings being asked for recognising correctness rather than their ability of drawing conclusions. We have made similar empirical research oriented towards the ability of recognising correctness of syllogistic reasoning.

We will report results of our investigations and some conclusions. They are particularly interesting from the point of view of the relation between logical competence and logical correctness.

[Johnson–Laird, Bara 1984] P. N. Johnson–Laird and B. G. Bara,Syllogistic inference, in Cognition 16 (1984), pp. 1–61.

Address: Institute of Philosophy, Warsaw University, Madalińskiego 42 m. 88, 02 – 544 Warszawa, Poland

E-mail: pbucholc@venus.ci.uw.edu.pl